OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals. 3/8x-1/2=1/8x +3/4

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }x-\frac{ 1 }{ 2}=\frac{ 1 }{ 8}x+\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I need to show all my work

OpenStudy (anonymous):

\[\frac{ 3x }{ 8 }-\frac{ 1x }{ 8 } = \frac{ 3 }{ 4 }+\frac{ 1 }{ 2 }\] taking lcm \[\frac{ 3x -x}{ 8 } = \frac{ 3 }{ 4 }+\frac{ 2 }{ 4 }\] simplifying \[\frac{ 2x}{ 8 } = \frac{ 3 +2}{ 4 }\] \[\frac{ x}{ 4 } = \frac{5}{ 4 }\] x = 4

OpenStudy (anonymous):

Sorry x = 5.